OpenStudy (anonymous):

Two numbers have a sum of 23 and a difference of 9. Find the two numbers

OpenStudy (anonymous):

your idea?

OpenStudy (anonymous):

none i have no idea how to come about this

OpenStudy (anonymous):

I give you trick, very easy, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

(sum + difference )/2 = big (sum - difference)/2 = small that's it

OpenStudy (anonymous):

.-.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

not at all

OpenStudy (anonymous):

you have 2 numbers need to find out, right?

OpenStudy (anonymous):

i found them out yes it is 16, and 7

OpenStudy (anonymous):

oh yea, you are so good.

OpenStudy (anonymous):

lol ok thx

OpenStudy (anonymous):

you see, sum =23 difference = 9 (sum + difference)/2 = (23+9)/2 = 32/2 =16 (the big number) (sum - difference )/2 =(23-9)/2 = 14/2 =7 (the small number) they are the answers, right? easy, right?